1) True or false?

            a.    n2 = O(n3)

            b.   2n2 + 1 = O(n2)

            c. n1/2 = O(log n)

            d.   log n = O(n1/2 )

            e.   log n + n1/2 = O(n2)

            f.    log n = O( n-1/2)

            g.    log n = O( 1/n )

            h.    log (n + 3) = q(n1/2)

            i.    n + n1/2= W( n2 - n )
Answer #1
a.    n2 = O(n3)
True

b.   2n2 + 1 = O(n2)
True

c. n1/2 = O(log n)
False

d.   log n = O(n1/2 )
True

e.   log n + n1/2 = O(n2)
True

f.    log n = O( n-1/2)
False

g.    log n = O( 1/n )
False

h.    log (n + 3) = W(n1/2)
False

i.    n + n1/2= W( n2 - n )
False